What is the probability that a radioactive atom having a mean life of 10 days decays during the fifth day?
(Ans.....0.39)

Grade:12

1 Answers

vikas askiitian expert
13 years ago

probablity = (1-e-(lamda)t)                         (lamda = decay constant)

    mean life = 1/lamda

      10 = 1/lamda

      lamda = 1/10 day-1

now at t = 5days

 probablity = (1-e-1/2)

                =(1- 1/e1/2)=0.39                     (value of e is around 2.8)